OpenStudy (anonymous):

Find the equation of a line tngent to F(x)=sq root of (2x+1) @ the point (4,3),Explain

OpenStudy (zehanz):

The slope of the tangent line at a point of a graph is equal to the derivative of the function. Here we have the function\[f(x)=\sqrt{2x+1}\]so we need to differentiate this function and then calculate f'(4). Now f is a composition of more than one function, so we need to use the Chain Rule. The derivative of \[\sqrt{x} \] is \[\frac{ 1 }{ 2\sqrt{x} }\] therefore \[f'(x)=\frac{ 1 }{ 2\sqrt{2x+1} }*2\] (the *2 is the derivative of 2x+1). This can be written as\[f'(x)=\frac{ 1 }{ \sqrt{2x+1} }\]Now calculate the derivative in x = 4 to find the slope of the tangent line:\[f'(4)=\frac{ 1 }{ \sqrt{2*4+1} }=\frac{ 1 }{ 3 }\] The equation of the tangent line is \[y=\frac{ 1 }{3 }x + b\]We now only need to find b. Because the point (4,3) is on this line, substituting 4 for x and 3 for y gives:\[b=y-\frac{ 1 }{ 3}x=3-\frac{ 1 }{ 3 }*4=\frac{ 8 }{ 3 }\]The equation is \[y=\frac{ 1 }{ 3 }x + \frac{ 8 }{ 3 }\]If you want to, this can also be written (by multiplying with 3 and arranging terms) as :\[x-3y+24=0\] Hope this helps! ZeHanz